Calculate the present value of the compound interest loan. (Round your answers to the nearest cent.)

$25,000 after 7 years at 4% if the interest is compounded in the following ways.

(a) annually
           $   

(b) quarterly
           $  
Answer #1

(a)

The present value is calculated below:

Present value \(=25,000 \times(1+0.04)^{-7}\)

Present value \(=18,997.95\)

(b)

Quarterly interest rate will be \((0.04 / 4)=0.01\)

The present value is calculated below:

Present value \(=25,000 \times(1+0.01)^{-28}\)

Present value \(=18,920.89\)